Two angles are complementary. They measure of the second angle is two degrees more than three times the measure of the first. Find the measures of the angles.
Answer by Mathtut(3670)

let the angles be x and y
:
y=3x+2.......eq 1
x+y=90.......eq 2
:
take y's value of 3x+2 from eq 1 and plug it into eq 2
:
x+3x+2=90
:
4x=88
:
x=22
:
y=3(22)+2=68